OpenStudy (yrivers36):

Write the equation of the line that passes through the points (-2,4) and (0,3). With explanation please.

OpenStudy (mimi_x3):

Use the gradient formula to find m m = -1/2 y = mx + b , (-2,4) 4 = -1/2(-2) + b 4 = 1 + b b = 3 y = -1/2x + 3

OpenStudy (yrivers36):

Thank you so much

OpenStudy (mimi_x3):

yw (: